OpenStudy (anonymous):

Calculate the number of years for an unvestment if $1000 to a double at an interest rate of 7.2% for each compounding period a) annually b) semi-annually c) monthly d) daily Please Show Full Solutions .. thank you

sam (.sam.):

after each period you get the ammount that you already had + 7.2% of that ammount. 1000 + 1000 (7.2/100) = 1000 + 1000 (72/1000) = 1000 + 72 = 1072. Next year you add 7.2% of 1072 to 1072. For semi-anual calculation, you can devide 7.2% by 2

OpenStudy (anonymous):

1.072 t = ln2 -> t = ln2/ 1.072 = .65 ( ~ 1 yr)

OpenStudy (anonymous):

2. 072 t = ln2 =>t = ln2/ 2.072 = .33 yr

OpenStudy (anonymous):

-> t = ln2/ 12.072 = .06 yr

OpenStudy (anonymous):

-> t = ln(2)/ 365.072 = .002 yr